Which sample contains the greatest number of atoms. A sample of I that contains 1.08×1024 atoms or a 8.32 mole sample of F?
Answer #1

Either convert one into mole or another into no of atoms to get the conversion

1 mole contains = 6.022 × 1023 atoms

Now converting 8.32 mole into no of atoms

8.32 mole contains = 8.32 × 6.022 × 1023 atoms

= 5.01 × 1024 atoms

So F contains more atoms